OpenStudy (anonymous):

What is the solution to the equation? 7x + 6 = –22 A. B. x = –4 C. D. x = 4

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@bohotness

OpenStudy (dumbsearch2):

Sure, I can help! :)

OpenStudy (anonymous):

thanks

OpenStudy (dumbsearch2):

From 7x + 6 = –22, minus 6 on both sides. Do that.

OpenStudy (dumbsearch2):

To take six out.

OpenStudy (dumbsearch2):

Refresh your page if you see weird symbols, I don't know why it does that.

OpenStudy (dumbsearch2):

We want to find X = "something".

OpenStudy (dumbsearch2):

So after you take six out, 7x = ?

OpenStudy (anonymous):

would X= 16?

OpenStudy (dumbsearch2):

No. Let's not jump there yet.

OpenStudy (dumbsearch2):

what is -22-6?

OpenStudy (anonymous):

-28 right

OpenStudy (dumbsearch2):

Right! So now that we took six out of "7x + 6 = –22", we have "7x + 0 = –28". That makes 7x + 0 = –28.

OpenStudy (anonymous):

do you need A C?

OpenStudy (dumbsearch2):

Why do you say do I need A C?

OpenStudy (anonymous):

there is no anwser for A C

OpenStudy (anonymous):

wait so the anwser is D

OpenStudy (dumbsearch2):

I never said the answer was A C

OpenStudy (anonymous):

oh ok

OpenStudy (dumbsearch2):

So now that we have 7x = -28, divide by 7 on both sides. what is -28/7? that makes -4. The answer is x = -4. This was a relatively simple linear equation.

OpenStudy (dumbsearch2):

Hope that helped, click "Best Response" because it did.

OpenStudy (anonymous):

would you help me with some more??

OpenStudy (dumbsearch2):

Yes, only if you put them in different questions though.

OpenStudy (dumbsearch2):

Make a new question.

OpenStudy (anonymous):

ok thanks

OpenStudy (dumbsearch2):

Then tag me if you want.